# Algebraic perturbation theory: traversable wormholes and generalized entropy beyond subleading order

### v1

@@ 3 Weighting of expectation values under a general perturbation
In this section, we compute the change in the entropy for a type II factor under a unitary perturbation of the original type III1 algebra. In particular, we compute the change in expectation values in the type II crossed product algebra, which become weighted by an exponential factor that depends on the perturbation. As we shall eventually apply this to a double-trace deformation of the thermofield double (TFD) state Gao:2016bin in sec. 4 below, we will have in mind that the unperturbed state $\Psi_{0}$ is the original TFD, while the perturbed state $\Psi$ is the TFD with the double-trace deformation turned on, but the formalism in this section is completely general, and applies to any two type III1 algebras related by some unitary transformation.
Before proceeding to the main computation in subsec. 3.4, we first clarify a potential confusion regarding how entropy can change under a unitary transformation in subsec. 3.1, and verify the corresponding change in the modular Hamiltonian in subsec. 3.2.
3.1 Semifinite partial traces and (non)unitary dynamics
Here, let us explicate what we mean by a change in entropy under a unitary transformation by returning to the more familiar type I algebra of quantum mechanics, where it is a basic fact such a transformation leaves the von Neumann entropy invariant. That is, if we have some density state $\rho$ that transforms as $\rho\mapsto U\rho U^{\dagger}$ under some unitary operator $U$, then the von Neumann entropy $S=-\mathrm{tr}\rho\ln\rho$ remains unchanged:
$S\mapsto-\mathrm{tr}\left[U\rho U^{\dagger}\ln(U\rho U^{\dagger})\right]=-\mathrm{tr}\left[U\rho U^{\dagger}U(\ln\rho)U^{\dagger}\right]=-\mathrm{tr}\rho\ln\rho\leavevmode\nobreak\ ,$ (12)
where the last step relied crucially on the cyclicity of the trace,
$\mathrm{tr}(AB)=\mathrm{tr}(BA)\leavevmode\nobreak\ .$ (13)
$\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}\leavevmode\nobreak\ .$ (14)
$\mathrm{tr}_{\bar{A}}\mathcal{O}=A_{1}A_{2}\,\mathrm{tr}\left(\bar{A}_{1}\bar{A}_{2}\right)=A_{1}A_{2}\,\mathrm{tr}\left(\bar{A}_{2}\bar{A}_{1}\right)\leavevmode\nobreak\ ,$ (15)
since the trace over the complement does not act on $\mathcal{H}_{A}$. [^6] If $[A_{1},A_{2}]=0$, then one obtains the same result as if the partial trace were cyclic on the full Hilbert space, but the latter is nonsensical since it does not even act on both factors. More pertinently, this means that if we consider a unitary transformation defined on the full Hilbert space, then
$\mathrm{tr}_{\bar{A}}\left(U\mathcal{O}U^{\dagger}\right)\neq\mathrm{tr}_{\bar{A}}\left(U^{\dagger}U\mathcal{O}\right)=\mathrm{tr}_{\bar{A}}\,\mathcal{O}\leavevmode\nobreak\ ,$ (16)
In other words, unitary evolution on $\mathcal{H}$ generally does not descend to unitary evolution on $\mathcal{H}_{A}$. One can of course still define a density state on the subsystem after unitary evolution on the whole,
$\rho^{\prime}_{A}=\mathrm{tr}_{\bar{A}}\left[U(\rho_{A}\otimes\rho_{\bar{A}})U^{\dagger}\right]\neq U\rho_{A}U^{\dagger}\leavevmode\nobreak\ ,$ (17)
$\rho^{\prime}_{A}=\sum_{i}E_{i}\rho_{A}E_{i}^{\dagger},$ (18)
Returning to our case of interest, it is impossible to define a partial trace on either the algebra or its commutant because both are type III1. For that reason, we invoke the crossed product construction to obtain a semifinite type II factor, which allows us to define a trace. As in the type I case just reviewed however, this trace will not be cyclic for a unitary operator acting on both the algebra and its commutant, so the global unitary transformation will still lead to non-trivial changes in density operators and expectation values defined on the subsystem. This is what allows the von Neumann entropy of the type II algebra (e.g., of the right exterior in the TFD) to change under the interaction Hamiltonian.
@@ 3.2 Covariant modular structure
Here, we establish some notation that will be used throughout the remainder of the paper, and in particular verify that the modular Hamiltonian transforms as expected under the perturbation. Let $\mathfrak{A}_{0}$ be a type III1 algebra acting on $\mathcal{H}_{0}$, and $\widehat{\mathfrak{A}}_{0}$, ^$\widehat{\mathcal{H}}_{0}$ the corresponding objects obtained by adjoining the modular charge $X_{0}$ as per the crossed product construction above. Similarly, let $\mathfrak{A}$ by the type III1 algebra algebra obtained from $\mathfrak{A}_{0}$ by applying a unitary transformation,
$\mathfrak{A}=U\mathfrak{A}_{0}U^{\dagger}\leavevmode\nobreak\ ,$ (19)
which acts on $\mathcal{H}$; the corresponding crossed product objects will be denoted $\widehat{\mathfrak{A}}$, $\widehat{\mathcal{H}}$. Since the unitary does not act on the adjoined $L^{2}(\mathbb{R})$ factor in the crossed product, it follows immediately that $\widehat{\mathfrak{A}}=U\widehat{\mathfrak{A}}_{0}U^{\dagger}$. Note that when applied to the TFD below, $\mathfrak{A}_{0}$ and $\mathfrak{A}$ will denote the full exterior algebra on both sides, whose corresponding type II algebras admit a decomposition of the form $\widehat{\mathfrak{A}}_{0}=\widehat{\mathfrak{A}}_{L,0}\otimes\widehat{\mathfrak{A}}_{R,0}$, and similarly for $\widehat{\mathfrak{A}}$, where the subscripts ,${}_{L},{}_{R}$ denote the left and right sides of the TFD (i.e., $\widehat{\mathfrak{A}}_{L,0}^{\prime}=\widehat{\mathfrak{A}}_{R,0}$).
@@
$\mathfrak{A}|{\Psi}\rangle=(U\mathfrak{A}_{0}U^{\dagger})U|{\Psi_{0}}\rangle=U\mathfrak{A}_{0}|{\Psi_{0}}\rangle\leavevmode\nobreak\ .$ (20)
Now, if we have access to density operators – which we do for the type II algebras $\widehat{\mathfrak{A}}_{0}$ and $\widehat{\mathfrak{A}}$ – then this suggests that $\widehat{\rho}_{0}=|{\widehat{\Psi}_{0}}\rangle\langle{\widehat{\Psi}_{0}}|$ also transforms to $\widehat{\rho}=U\widehat{\rho}_{0}U^{\dagger}$. [^7] Note that this does not contradict (17) because here, $\rho$ is the density operator on the full system, not one obtained via a partial trace. Formally, if we think of $\rho=e^{-H_{\mathrm{mod}}}$, then this suggests that the modular operator transforms as
$H_{\mathrm{mod}}=UH_{\mathrm{mod},0}U^{\dagger}\leavevmode\nobreak\ .$ (21)
That this preserves the separating property of the vacuum state again follows immediately from the unperturbed state:
$H_{\mathrm{mod}}|{\Psi}\rangle=(UH_{\mathrm{mod},0}U^{\dagger})U|{\Psi_{0}}\rangle=UH_{\mathrm{mod},0}|{\Psi_{0}}\rangle=0\leavevmode\nobreak\ .$ (22)
@@
${\Psi}\right|{M(t)aM(-t)b}\left|{\Psi}\right>=\left<{\Psi}\right|{bM(t+i\beta)aM(-t-i\beta)}\left|{\Psi}\right>\leavevmode\nobreak\ ,\qquad\forall a,b\in\mathfrak{A}\leavevmode\nobreak\ .$ (23)
@@
${\Psi}\right|{aM(-t)b}\left|{\Psi}\right>=\left<{\Psi}\right|{bM(t+i\beta)a}\left|{\Psi}\right>\leavevmode\nobreak\ .$ (24)
@@
${\Psi_{0}}\right|{U^{\dagger}aUM_{0}(-t)U^{\dagger}bU}\left|{\Psi_{0}}\right>$ =${\Psi_{0}}\right|{U^{\dagger}bUM_{0}(t+i\beta)U^{\dagger}aU}\left|{\Psi_{0}}\right>$ (25)
${\Psi_{0}}\right|{a_{0}M_{0}(-t)b_{0}}\left|{\Psi_{0}}\right>$ =${\Psi_{0}}\right|{b_{0}M_{0}(t+i\beta)a_{0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ ,$
@@
We note in passing that we may also define a modular conjugation $J$ for the new algebras in terms of the original one, $J_{0}$, by acting with this same unitary as $J=UJ_{0}U^{\dagger}$. Starting from an element $a_{R}\in\mathfrak{A}_{R}$, we then have
$Ja_{R}J=UJ_{0}U^{\dagger}a_{R}UJ_{0}U^{\dagger}=UJ_{0}\,a_{R,0}\,J_{0}U^{\dagger}=Ua_{L,0}U^{\dagger}=a_{L}\in\mathfrak{A}_{L}\leavevmode\nobreak\ .$ (26)
@@ 3.3 General change in weighting of expectation values
Preliminaries thus dispensed with, we can now proceed to our main interest: the change in the weighting of expectation values in the type II theory under a unitary perturbation. As reviewed above, the type III systems $\mathfrak{A}_{0}$, $\mathfrak{A}$ do not admit a tensor product factorization, but the type II systems $\widehat{\mathfrak{A}}_{0}$, $\widehat{\mathfrak{A}}$ do. Accordingly, suppose we factorize these algebras into left ($\widehat{\rho}_{L,0}$, $\widehat{\rho}_{L}$) and right ($\widehat{\rho}_{R,0}$, $\widehat{\rho}_{R}$) density matrices. In the application to the traversable wormhole below for example, $\widehat{\rho}_{R,0}$ and $\widehat{\rho}_{R}$ will be the density matrices of the right exterior (that is, outside the black hole) algebra before and after the deformation, respectively. While the transformation (19) is unitary on the full algebra, it does not necessarily act unitarily on either of these factors as explained above, which allows for the possibility of a change in the von Neumann entropy. Since this can only be rigorously defined in the type II theory, we wish to compute
$\Delta S=S(\widehat{\rho}_{R})-S(\widehat{\rho}_{R,0})\leavevmode\nobreak\ ,$ (27)
with the von Neumann entropies on the right-hand side given by (10), with $K\in\{\widehat{\rho}_{R},\widehat{\rho}_{R,0}\}$.
However, doing this requires that we make more precise some notational liberties taken in the previous crossed product works reviewed above. In particular, note that in (7), the expectation value ${\Psi}\right|{\widehat{a}}\left|{\Psi}\right>$ is an operator-valued function of the modular charge $X$. But as an operator, it is not meaningful to integrate this over the reals as written. Intuitively, what is meant is that we integrate over the (continuous) spectrum of $X$, whose domain is $\mathbb{R}$. This is relevant here because, as we will see, the weighting from the perturbation results in a factor containing commutators of operators which is non-diagonal in the energy eigenbasis, so it is helpful to make this slightly cavalier step to c-numbers more explicit. To that end, observe that when we write
@@
$g^{1/2}(\lambda)=\langle\lambda|g^{1/2}\rangle\leavevmode\nobreak\ ,$ (29)
@@
$\int\!\mathrm{d}X\,g(X)\coloneqq\int\!\mathrm{d}\lambda\,|\langle\lambda|g^{1/2}\rangle|^{2}=1\leavevmode\nobreak\ ,$ (30)
where we define the non-rigorous expression on the far left-hand side by the expression in the middle. The physical interpretation of this is that $|\langle\lambda|g^{1/2}\rangle|^{2}$ is the probability of the energy eigenstate $\lambda$ given the prepared state $g^{1/2}$. A correlation function in $L^{2}(\mathbb{R})$ is similarly computed by inserting a complete basis of these eigenstates. In particular, given an arbitrary function $f(X)$,
$f(X)\right>_{L^{2}(\mathbb{R})}=\langle{g^{1/2}}|f(X)|{g^{1/2}}\rangle=\int\!\mathrm{d}\lambda\langle{g^{1/2}}|f(X)|{\lambda}\rangle\langle\lambda|g^{1/2}\rangle=\int\!\mathrm{d}\lambda\,f(\lambda)|\langle\lambda|g^{1/2}\rangle|^{2}\leavevmode\nobreak\ ,$ (31)
@@
$\mathrm{tr}\,\widehat{a}$ =${\widehat{\Psi}}\right|{\widehat{a}K^{-1}}\left|{\widehat{\Psi}}\right>$ (32)
@@
$=\int\!\mathrm{d}\lambda\,e^{\lambda}\left<{\Psi}\right|{\widehat{a}}\left|{\Psi}\right>\leavevmode\nobreak\ ,$
@@
Relating this to the undeformed expectation value is then quite straightforward: we have
$|{\Psi}\rangle=U|{\Psi_{0}}\rangle\leavevmode\nobreak\ ,\qquad\widehat{a}=U\widehat{a}_{0}U^{\dagger}\leavevmode\nobreak\ ,\qquad K^{-1}=UK^{-1}_{0}U^{\dagger}\leavevmode\nobreak\ ,$ (33)
where the last follows from the assumption below (6) that $g^{-1}$ is analytic. Crucially, note again that since $U$ does not act on $L^{2}(\mathbb{R})$, it commutes with any function of $X$. Hence, inserting the relations (33) in the third line in (32), all factors of $U$, $U^{\dagger}$ commute through and combine to $\mathbb{1}$, so that the last line becomes
@@
where ${\Psi_{0}}\right|{\widehat{a}_{0}}\left|{\Psi_{0}}\right>$ as well as $\lambda(\lambda_{0})$ is a function of $\lambda_{0}\neq\lambda$. Since the spectral measure is unique, [^9] A unitary transformation preserves eigenvalues and thus must preserve the spectral measure. the Jacobian $\mathrm{d}\lambda/\mathrm{d}\lambda_{0}=1$, so this becomes simply
$\mathrm{tr}\,\widehat{a}=\int\!\mathrm{d}\lambda_{0}\,e^{\lambda(\lambda_{0})}\left<{\Psi_{0}}\right|{\widehat{a}_{0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ .$ (35)
In other words, the expectation value of observables in the perturbed type II algebra is the same as that in the unperturbed algebra up to a weighting factor $e^{\lambda(\lambda_{0})}$. Intuitively, this is the expectation value of $Ue^{X_{0}}U^{\dagger}$ in the energy eigenbasis, where $X_{0}$ denotes the modular charge in the unperturbed theory. To see this, suppose we work directly with the non-rigorous notation in (7), and apply the unitary transformation in (33) to the modular charge (i.e., the one-sided Hamiltonian), i.e., $X=UX_{0}U^{\dagger}$: [^10] Strictly speaking, since $X$, $X_{0}$ act on the $L^{2}(\mathbb{R})$ factor and not on the TFD Hilbert space, we should denote this by some other operator $\widetilde{U}$. This technicality will become clear shortly.
@@
$=\int\!\mathrm{d}X_{0}\,\left(Ue^{X_{0}}U^{\dagger}\right)\langle{\Psi_{0}}|\widehat{a}_{0}|{\Psi_{0}}\rangle\leavevmode\nobreak\ ,$
where in the last step, we used the fact that a unitary transformation does not alter the spectral measure, and the freedom to move unitary operators in/out of exponentials at will. Since both $U$ and $X_{0}$ are operators however, these must be evaluated in some state [^11] As noted in Witten:2021unn , it does not matter which state we choose, since the factor $g$ will drop out of the final expression. This reflects the fact that the spectrum ultimately depends on the operator, not the state. before the integral over the spectrum can be performed, cf. (30).
Let us make the connection to the weight $e^{\lambda(\lambda_{0})}$ in (34) more explicit. In our application below, we will be interested in the case in which $X$ is identified with the suitably-normalized left modular charge $\widetilde{H}_{L}$, and is related to the unperturbed charge by [^12] The tilde notation concerns the $1/N$ subtleties postponed above, and distinguishes the normalized ${\widetilde{H}_{L}=H_{L}/N}$ vs. unnormalized modular charge $H_{L}$; this will be discussed in subsec. 4.2, and the distinction will be relevant momentarily.
$\widetilde{H}_{L}=\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}\leavevmode\nobreak\ ,$ (37)
@@
$Ue^{\widetilde{H}_{L,0}}U^{\dagger}=e^{\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}}\leavevmode\nobreak\ ,$ (38)

### v2

@@ 3 Weighting of expectation values under a general perturbation
In this section, we compute the entropy for a type II factor following a unitary perturbation of the original system. In particular, we compute the effect on expectation values in the type II crossed product algebra, which become weighted by an exponential factor that depends on the perturbation. As we shall eventually apply this to a double-trace deformation of the thermofield double (TFD) state Gao:2016bin in sec. 4 below, we will have in mind that the unperturbed state $\Psi_{0}$ is the original TFD, while the perturbed state $\Psi$ is the TFD with the double-trace deformation turned on, but the formalism in this section is completely general, and applies to any two type III1 algebras related by some unitary transformation.
We first verify the covariance of the modular structure under a unitary transformation in subsec. 3.1. Then in subsec. 3.2, we analyze the change in expectation values – in particular that appearing in the trace formula for type II algebras – under the perturbation. This mathematical result is then applied to the von Neumann entropy in subsec. 3.3. For pedagogical purposes, we have included a contextual discussion of how the entropy of a subsystem changes under a unitary transformation in appendix A.
@@ 3.1 Covariant modular structure
Here, we establish some notation that will be used throughout the remainder of the paper, and in particular verify that the modular Hamiltonian transforms as expected under the perturbation. Let $\mathfrak{A}_{0}$ be a type III1 algebra acting on $\mathcal{H}_{0}$, and $\widehat{\mathfrak{A}}_{0}$, $\widehat{\mathcal{H}}_{0}$ the corresponding objects obtained by adjoining the modular charge $X_{0}$ as per the crossed product construction above. Similarly, let $\mathfrak{A}$ be the type III1 algebra algebra obtained from $\mathfrak{A}_{0}$ by applying a unitary transformation,
$\mathfrak{A}=U\mathfrak{A}_{0}U^{\dagger}\penalty 10000\ ,$ (12)
which acts on $\mathcal{H}$; the corresponding crossed product objects will be denoted $\widehat{\mathfrak{A}}$, $\widehat{\mathcal{H}}$. Since the unitary does not act on the adjoined $L^{2}(\mathbb{R})$ factor in the crossed product, it follows immediately that $\widehat{\mathfrak{A}}=U\widehat{\mathfrak{A}}_{0}U^{\dagger}$. Note that when applied to the TFD below, $\mathfrak{A}_{0}$ and $\mathfrak{A}$ will denote the full exterior algebra on both sides, whose corresponding type II algebras admit a decomposition of the form $\widehat{\mathfrak{A}}_{0}=\widehat{\mathfrak{A}}_{L,0}\otimes\widehat{\mathfrak{A}}_{R,0}$, and similarly for $\widehat{\mathfrak{A}}$, where the subscripts $_{L},{}_{R}$ denote the left and right sides of the TFD (i.e., $\widehat{\mathfrak{A}}_{L,0}^{\prime}=\widehat{\mathfrak{A}}_{R,0}$). We emphasize that the one-sided subalgebras are not unitarily related, i.e., $\widehat{\mathfrak{A}}_{R}\neq U\widehat{\mathfrak{A}}_{R,0}U^{\dagger}$, since a generic unitary $U$ will mix operators between the left and right.
@@
$\mathfrak{A}|{\Psi}\rangle=(U\mathfrak{A}_{0}U^{\dagger})U|{\Psi_{0}}\rangle=U\mathfrak{A}_{0}|{\Psi_{0}}\rangle\penalty 10000\ .$ (13)
Now, if we have access to density operators – which we do for the type II algebras $\widehat{\mathfrak{A}}_{0}$ and $\widehat{\mathfrak{A}}$ – then this suggests that $\widehat{\rho}_{0}=|{\widehat{\Psi}_{0}}\rangle\langle{\widehat{\Psi}_{0}}|$ also transforms to $\widehat{\rho}=U\widehat{\rho}_{0}U^{\dagger}$. Formally, if we think of $\rho=e^{-H_{\mathrm{mod}}}$, then this suggests that the modular Hamiltonian transforms as
$H_{\mathrm{mod}}=UH_{\mathrm{mod},0}U^{\dagger}\penalty 10000\ .$ (14)
That this preserves the annihilation effect of $H_{\rm mod}$ on the perturbed vacuum state again follows immediately from that of the unperturbed state:
$H_{\mathrm{mod}}|{\Psi}\rangle=(UH_{\mathrm{mod},0}U^{\dagger})U|{\Psi_{0}}\rangle=UH_{\mathrm{mod},0}|{\Psi_{0}}\rangle=0\penalty 10000\ .$ (15)
@@
$\left<{\Psi}\right|{M(t)aM(-t)b}\left|{\Psi}\right>=\left<{\Psi}\right|{bM(t+i\beta)aM(-t-i\beta)}\left|{\Psi}\right>\penalty 10000\ ,\qquad\forall a,b\in\mathfrak{A}\penalty 10000\ .$ (16)
@@
[…]
@@
$\left<{\Psi_{0}}\right|{U^{\dagger}aUM_{0}(-t)U^{\dagger}bU}\left|{\Psi_{0}}\right$> $=\left<{\Psi_{0}}\right|{U^{\dagger}bUM_{0}(t+i\beta)U^{\dagger}aU}\left|{\Psi_{0}}\right$> (18)
[…]
@@
We note in passing that we may also define a modular conjugation $J$ for the new algebras in terms of the original one, $J_{0}$, by acting with this same unitary as $J=UJ_{0}U^{\dagger}$. Starting from an element $a_{R}\in\mathfrak{A}_{R}$, we then have [^9] Note that for simplicity, we have here assumed that elements do not mix between left and right subalgebras under the unitary, but the relation holds even if this is relaxed. That is, suppose that instead we chose an element $a_{R}$ such that $U^{\dagger}a_{R}U=a_{L,0}$. Then, starting from the middle expression, we have $UJ_{0}a_{L,0}J_{0}U^{\dagger}=Ua_{R,0}U^{\dagger}=a_{L}$, where the last step follows from the symmetry of the (unitary action on the) system.
$Ja_{R}J=UJ_{0}U^{\dagger}a_{R}UJ_{0}U^{\dagger}=UJ_{0}\,a_{R,0}\,J_{0}U^{\dagger}=Ua_{L,0}U^{\dagger}=a_{L}\in\mathfrak{A}_{L}\penalty 10000\ .$ (19)
@@ 3.2 General change in weighting of expectation values
We can now proceed to our main interest: the change in the weighting of expectation values in the type II theory under a unitary perturbation. As reviewed above, the type III systems $\mathfrak{A}_{0}$, $\mathfrak{A}$ do not admit a tensor product factorization, but the type II systems $\widehat{\mathfrak{A}}_{0}$, $\widehat{\mathfrak{A}}$ do. Accordingly, suppose we factorize these algebras into left ($\widehat{\rho}_{L,0}$, $\widehat{\rho}_{L}$) and right ($\widehat{\rho}_{R,0}$, $\widehat{\rho}_{R}$) density matrices. In the application to the traversable wormhole below for example, $\widehat{\rho}_{R,0}$ and $\widehat{\rho}_{R}$ will be the density matrices of the right exterior (that is, outside the black hole) algebra before and after the deformation, respectively. While the transformation (12) is unitary on the full algebra of both sides, it does not necessarily act unitarily on either of these factors as explained above (see also appendix A for a pedagogical treatment). This is exactly the quantum information-theoretic explanation of the Stinespring dilation theorem, which always allows us to pass to a larger Hilbert space to turn a quantum channel into a unitary operator. Thus, in general, one expects a change in the entropy following a unitary operation on the full system. Concretely, we wish to compute $S(\widehat{\rho}_{R})$, the von Neumann entropy given by (10), with $K=\widehat{\rho}_{R}$.
However, doing this requires that we make more precise some notational liberties taken in the previous crossed product works reviewed above. In particular, note that in (7), the expectation value $\left<{\Psi}\right|{\widehat{a}}\left|{\Psi}\right$> is an operator-valued function of the modular charge $X$. But as an operator, it is not meaningful to integrate this over the reals as written. Intuitively, what is meant is that we integrate over the (continuous) spectrum of $X$, whose domain is $\mathbb{R}$. This is relevant here because, as we will see, the weighting from the perturbation results in a factor containing commutators of operators which is non-diagonal in the energy eigenbasis, so it is helpful to make this slightly cavalier step to c-numbers more explicit. To that end, observe that when we write
@@
$g^{1/2}(\lambda)=\langle\lambda|g^{1/2}\rangle\penalty 10000\ ,$ (21)
@@
$\int\!\mathrm{d}X\,g(X)\coloneqq\int\!\mathrm{d}\lambda\,|\langle\lambda|g^{1/2}\rangle|^{2}=1\penalty 10000\ ,$ (22)
where we define the formal expression on the far left-hand side by the expression in the middle. The physical interpretation of this is that $|\langle\lambda|g^{1/2}\rangle|^{2}$ is the probability of the energy eigenstate $\lambda$ given the prepared state $g^{1/2}$. A correlation function in $L^{2}(\mathbb{R})$ is similarly computed by inserting a complete basis of these eigenstates. In particular, given an arbitrary function $f(X)$,
$\left<f(X)\right>_{L^{2}(\mathbb{R})}=\langle{g^{1/2}}|f(X)|{g^{1/2}}\rangle=\int\!\mathrm{d}\lambda\langle{g^{1/2}}|f(X)|{\lambda}\rangle\langle\lambda|g^{1/2}\rangle=\int\!\mathrm{d}\lambda\,f(\lambda)|\langle\lambda|g^{1/2}\rangle|^{2}\penalty 10000\ ,$ (23)
@@
$\mathrm{tr}\,\widehat{a}$ $=\left<{\widehat{\Psi}}\right|{\widehat{a}K^{-1}}\left|{\widehat{\Psi}}\right$> (24)
@@
[…]
@@
If the one-sided algebras were unitarily related, then relating this to the undeformed expectation value would be quite straightforward, since then we have
$|{\Psi}\rangle=U|{\Psi_{0}}\rangle\penalty 10000\ ,\qquad\qquad\widehat{a}=U\widehat{a}_{0}U^{\dagger}\penalty 10000\ .$ (25)
However, as emphasized above, $U$ is only unitary when acting on the full two-sided algebra; hence, the second expression may fail, since for some $\widehat{a}_{0}\in\widehat{\mathfrak{A}}_{R,0}$, the unitary may map it to $\widehat{a}\in\widehat{\mathfrak{A}}_{L}$. If we wish to remain strictly within the same algebra, then the relation between $\widehat{a}_{0}$ and $\widehat{a}$ must be modeled as a quantum channel. Schematically, we can relate $\widehat{a}$ to $\widehat{a}_{0}$ using the Kraus operators $V_{i}$, for $i\in\mathcal{I}$ where $\mathcal{I}$ is an appropriate index set of the quantum channel obtained from the deformation,
$\widehat{a}=\sum_{i}V_{i}\widehat{a}_{0}V_{i}^{\dagger}.$ (26)
The trace in the generic case would then be computed via
$\mathrm{tr}\,\widehat{a}=\int\!\mathrm{d}\lambda\,e^{\lambda}\sum_{i}\langle{\Psi_{0}}|U^{\dagger}V_{i}\widehat{a}_{0}V_{i}^{\dagger}U|{\Psi}\rangle\penalty 10000\ .$ (27)
Here, we will limit ourselves to the subset of operators that do not mix under the unitary, so that (25) holds, postponing the more general relation by quantum channels to future work. In that case, the Kraus decomposition simply collapses to $\widehat{a}=U\widehat{a}_{0}U^{\dagger}$. All factors of $U$, $U^{\dagger}$ commute through and combine to $\mathbb{1}$, so that the expression simplifies
@@
where $\left<{\Psi_{0}}\right|{\widehat{a}_{0}}\left|{\Psi_{0}}\right$> as well as $\lambda(\lambda_{0})$ is a function of $\lambda_{0}\neq\lambda$. Importantly, the Jacobian will be non-trivial, since the spectral measure is not preserved under the action of $U$ (since this is not a unitary on the one-sided algebra); we will return to this at the end of the present subsection.
In other words, the expectation value of observables in the perturbed type II algebra is the same as that in the unperturbed algebra up to a weighting factor $e^{\lambda(\lambda_{0})}$ times the Jacobian. Intuitively, the former is the expectation value of $Ue^{X_{0}}U^{\dagger}$ in the energy eigenbasis, where $X_{0}$ denotes the modular charge in the unperturbed theory. To see this, suppose we work directly with the purely formal notation in (7), and apply the unitary transformation in (25) to the modular charge (i.e., the one-sided Hamiltonian), i.e., $X=UX_{0}U^{\dagger}$: [^11] This is a slight abuse of notation, since $X$ and $X_{0}$ act only on the $L^{2}(\mathbb{R})$ factors (cf. the comment about $U$ commuting with $X$ above). Properly, what we mean is that we act on the QFT with $U$, i.e., $UH_{L,0}U^{\dagger}$ (formally) to obtain $H_{L}$, and then normalize appropriately to obtain $X=X_{0}+\delta X_{0}$; this technicality will be discussed shortly.
@@
$=\int\!\mathrm{d}(UX_{0}U^{\dagger})\,\left(Ue^{X_{0}}U^{\dagger}\right)\langle{\Psi_{0}}|\widehat{a}_{0}|{\Psi_{0}}\rangle\penalty 10000\ .$
Since both $U$ and $X_{0}$ are operators however, these must be evaluated in some state [^12] As noted in Witten:2021unn , it does not matter which state we choose, since the factor $g$ will drop out of the final expression. This reflects the fact that the spectrum ultimately depends on the operator, not the state. before the integral over the spectrum can be performed, cf. (22). Importantly, note that this relation implies that the deformed algebra inherits the constant ambiguity in the von Neumann entropy from the original algebra, cf. footnote 8. This constant arises from the freedom to shift $X_{0}\mapsto X_{0}+c$, and hence is unaffected by the transformation, so that the ambiguity in the definition of the trace is the same in both algebras. We will comment further on this below.
Let us make the connection to the weight $e^{\lambda(\lambda_{0})}$ in (28) more explicit. In our application below, we will be interested in the case in which $X$ is identified with the suitably-normalized left modular charge $\widetilde{H}_{L}$, and is related to the unperturbed charge by [^13] The tilde notation concerns the $1/N$ subtleties postponed above, and distinguishes the normalized $\widetilde{H}_{L}=(H_{L}-\langle H\rangle_{\Psi})/N$} vs. unnormalized modular charge $H_{L}$; this will be discussed in subsec. 4.2, and the distinction will be relevant momentarily.
$\widetilde{H}_{L}=\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}\penalty 10000\ ,$ (30)
@@
$Ue^{\widetilde{H}_{L,0}}U^{\dagger}=e^{\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}}\penalty 10000\ ,$ (31)